How do I solve the linear system 3x - 4y = 1 and y = - 2x + 8 and write my anwser for the y coordinate?
Answer by rfer(16322)

3x-4(-2x+8)=1
3x+8x-32=1
11x=33
x=3
-----------
3*3-4y=1
9-4y=1
-4y=-8
y=2
-------------
(3,2)
----------------
proof
2=-2*3+8